# Foundation Models in Electrocardiogram: A Review

### v2

@@ 4 Methodology Perspective
4.1 Architecture Design
As shown in Figure 2, we first investigate the architecture of ECG-FMs, including backbone models and Transformer structure, concentrating on the underlying mechanisms that determine their capabilities and the applications across different ECG-based scenarios.
3.1.1 Transformer-based Models.
The architecture of ECG-FMs has increasingly embraced the Transformer model, which was originally developed for NLP tasks [119]. The core feature of the Transformer lies in its use of the attention mechanism, enabling the model to dynamically process different parts of the input data in parallel. The attention mechanism maps a query and a set of key-value pairs to an output, which is a weighted sum of the values weighted by the similarity between the query and each key. Since the Transformer does not rely on any convolution or recurrence, positional encoding is employed to represent the relative positions of input tokens, facilitating the capture of long-range dependencies in sequences. The scalability and efficacy of the Transformer in discerning intricate data patterns have resulted in its widespread adoption outside of NLP, such as in CV, audio, medicine, and beyond.
The original Transformer model utilizes an encoder-decoder architecture with a stack of multi-head self-attention blocks. However, the most suitable transformer mode for ECG analysis is still a subject of debate, with options including encoder-only, decoder-only, and encoder-decoder [49]. Each mode offers distinct properties and is tailored to specific scenarios. The encoder-only structure, commonly found in ViTs, BERTs, and CLIPs, processes the entire input ECG sequence at once and predicts based on the understanding of the full context, making it ideal for tasks that require interpreting the entire input data, such as cardiac disease detection [130]. On the other hand, the decoder-only mode, typically seen in GPTs, LLaMAs, and DALL-Es, showcases outstanding generalization capabilities during large-scale ECG pre-training. Its proficiency in tasks such as zero-shot inference [67, 135], ECG synthesis [16], and ECG chatbot assistant [69, 136] highlights its versatility in application. Additionally, prominent research combines an encoder-only model and a decoder-only model to create a sophisticated encoder-decoder structure, achieving strong performance in diverse tasks [50, 118, 136], besides using a simple Transformer [31, 143] in pre-training. The variety of architectural choices underscores the potential and necessity for further exploration in this field.
The Transformer-based backbone of pre-trained ECG-FMs typically includes six categories: ViT-series, BERT-series, CLIP-series, GPT-series, LLaMA-series, and DAll-E-series.
  • ViT-series models. ViT was originally developed to revolutionize image recognition by directly applying the Transformer architecture to image patches [21]. Unlike CNNs that rely on local feature extraction through hierarchical layers, ViT takes a unique approach by linearly embedding images into non-overlapping vectors and treating them as input tokens for the Transformer encoder. The multi-head self-attention module in ViT empowers it to capture long-range dependencies across the entire image, setting a new standard for image analysis. ViT comes in two main variants: ViT-Base and ViT-Large. ViT-Large, with its 24 layers, 1024 hidden dimensions per layer, and 16 self-attention heads, represents a leap in architectural complexity. This deeper architecture equips ViT-Large with unparalleled capacity to capture complex patterns and long-range dependencies, making it particularly effective for handling large-scale datasets. In the realm of ECG-FM, ViT has emerged as a pre-trained model, enabling the extraction of spatial and temporal cardiac features from ECG plots or time-frequency images [87, 138]. Furthermore, ViT has also been leveraged in several studies to process ECG sequences by segmenting signals into non-overlapping time series windows [44, 103, 1, 131].
  • BERT-series models. BERT [19], a pre-trained bidirectional Transformer model, leverages contextual information from unlabeled data through masked language modeling (MLM). Its pre-training stage involves two unsupervised tasks: MLM and next sentence prediction (NSP). In the MLM task, approximately 15% of the input tokens are randomly masked with a special [MASK] token, prompting the model to predict the masked part using the surrounding context. The NSP task enables the model to comprehend sentence-level coherence by determining whether the sentence label is the ”actual next”. During the fine-tuning stage, BERT encodes a concatenated text pair using a bidirectional cross-attention mechanism between two sentences for downstream tasks. Expanding on BERT’s success in NLP, numerous BERT variants have emerged in other domains, such as BEiT [8] in CV, and BioBERT [61], ClinicalBERT [5] and BioLinkBERT [134] in medicine.
    BEiT leverages a discrete variational autoencoder (dVAE) to tokenize an image into patches and discrete visual tokens, subsequently replacing randomly masked image patches with a special mask embedding. These are then fed into a ViT to learn long-range dependencies. The pre-training task of BEiT, masked image modeling (MIM), closely resembles BERT’s MLM task, aiming to predict the visual tokens of the original image based on the encoding vectors of the corrupted image. Additionally, task-specific layers are directly appended to the pre-trained encoder to fine-tune the parameters on downstream tasks. In the realm of ECG-FM, ECG-GPT [50] adopts BEiT pre-trained on ImageNet dataset as the encoder to capture the complex visual features in ECG images. HeartBEiT [118] leverages BEiT in the masked image modeling phase to learn contextual information from tokenized representations.
    BioBERT represents a groundbreaking contextual model that explores the potential of BERT within the clinical and biomedical domains [61]. It involves pretraining a BERT model using a corpus of biomedical research articles sourced from PubMed abstracts and PMC full-text articles. BioBERT utilizes the WordPiece tokenization technique [127] to represent new words with frequent subwords, effectively addressing the out-of-vocabulary issue. Requiring minimal task-specific architectural modification, BioBERT surpasses BERT across various biomedical text mining tasks, including named entity recognition, relation extraction, and question answering.
    ClinicalBERT is a specialized version of the BERT model fine-tuned on clinical texts to address the unique linguistic characteristics of medical narratives such as physician notes and discharge summaries. Unlike BioBERT, it demonstrates superior performance in named entity recognition and medical natural language inference by introducing clinical notes in the pre-training stage [5]. However, it may not be as effective on de-identification tasks, where BioBERT might outperform it due to the nature of how protected health information is masked in clinical text. ClinicalBERT is also widely used in the field of ECG analysis. For example, ETP [66] and METS [63] employ ClinicalBERT, pre-trained on clinical notes and bio-clinical articles, as the text-encoding branch of the CLIP framework.
    BioLinkBERT is a pre-trained LinkBERT [134] specifically tailored for the biomedical domain, leveraging the citation links within PubMed articles. In contrast to traditional clinical BERT approaches, BioLinkBERT treats the corpus as a dynamic graph of interconnected documents, utilizing citation links to enhance the learning context. Through a dual focus on MLM and document relation prediction (DRP), BioLinkBERT captures intricate dependencies and relationships across documents, leading to superior performance in tasks requiring multi-hop reasoning and comprehension of complex inter-document relations. For example, CQA-ESI [136] utilizes BioLinkBERT as the textual encoder, which effectively embeds medical terminologies owing to its efficacy in processing biomedical text data.
  • CLIP-series models. CLIP has gained prominence as a multimodal model with zero-shot capabilities, excelling in various classification benchmarks using natural language instructions, without the need for direct optimization for each specific benchmark [98]. Trained on the extensive WebImageText [141] dataset, which consists of about 400 million pairs of images and corresponding text collected from the internet, CLIP significantly reduces reliance on expensive labeled datasets while enabling broader generalization. It utilizes separate Transformer encoders for text and image inputs, with ViT serving as the visual encoder and GPT-2 Transformer as the text encoder. By concurrently training an image and a text encoder, CLIP maximizes the similarity between embeddings of paired image-text samples while minimizing similarity with other combinations, establishing robust connections between visual representations and language text. This unique design facilitates seamless zero-shot transfer to a wide range of downstream tasks. Several following works have been proposed to enhance the pre-training performance of CLIP and customize it for different domains. For instance, DeCLIP [142], SLIP [86], and CYCLIP [33] leverage additional uni/multi-modality supervision to strengthen the model’s capabilities. Additionally, FLIP [133] introduces fine-grained alignment to the model.
    Despite the effectiveness of these CLIP variants on large-scale datasets, they encounter challenges when applied to medical image-text datasets due to their smaller scale and the potential for misclassification caused by individual differences among patients with the same disease. To tackle these challenges, practitioners have developed specialized CLIP models for the medical domain. MedCLIP [124] decouples images and texts and replaces the InfoNCE loss with a semantic matching loss based on medical knowledge to address false negatives. BiomedCLIP [139] replaces the GPT-2 Transformer with a pre-trained PubMedBERT [35] as the text encoder, showcasing exceptional performance in cross-modal retrieval, zero-shot image classification, and medical visual question answering.
  • GPT-series models. GPT is an autoregressive language model that generates human-like text using the Transformer architecture, initially introduced by [96] for natural language understanding tasks. The parameters learned during unsupervised pre-training are fine-tuned for specific tasks using annotated datasets. Trained on over 40 GB of text from public websites, GPT-2 [98] has demonstrated its capability to produce coherent and contextually relevant paragraphs. However, its practical application was limited due to performance issues. To address this, a more robust model, GPT-3 [13], was developed. GPT-3 employs a similar architecture to GPT-2 but implements alternating dense and locally banded sparse attention patterns in the transformer layers. With 175 billion parameters, GPT-3 can perform tasks without extensive retraining and significantly outperforms its predecessors in zero-shot and few-shot learning scenarios. The latest iteration, GPT-4 [2], expands the input prompt to include multimodal inputs such as image and audio, addressing a limitation of GPT-3, which only accepts text inputs. In the field of ECG analysis, several studies have utilized frozen GPT models to generate potential waveform details from medical knowledge bases [136, 135]. Other practitioners have attempted to create ECG-oriented GPT models, such as ECG-GPT [50] and SignalGPT [69]. To mitigate diagnostic bias and improve the effectiveness of GPT fine-tuning, these specialized models first generate objective descriptions by analyzing the input multimodal ECG data, and then re-analyze the generated text using prompts resembling physicians’ corrections.
  • LLaMA-series models. LLaMA represents a significant leap forward in language generation, which is trained on a vast dataset of over 1.4 trillion tokens and enhanced with key modifications to existing pre-trained models. Apart from GPT-based LLMs, LLaMA employs innovative techniques for stability and performance, including normalizing at the input of each Transformer sub-layer, replacing the ReLU activation function with the SwiGLU [105] activation function, and utilizing rotary positional embeddings (RoPE) [112] instead of the absolute positional embeddings [15]. Furthermore, LLaMA offers specialized variants such as LLaMA-2 [117] and LLaMA-3 [22], tailored for dialogue applications, surpassing open-source chat models on various benchmarks. LLaMA-2 introduces enhanced context length and grouped-query attention (GQA) for faster inference, while LLaMA-3 optimizes the RoPE base frequency and tokenizer compression rates to reduce training compute dependency. Beyond its prowess in NLP, LLaMA-based models demonstrate exceptional performance in zero-shot ECG classification and sleep apnea detection [135].
  • DAll-E-series models. DALL-E [101] is a cutting-edge language model that employs transformer-based technology to generate images from text descriptions through a two-stage training procedure. In the initial stage, a dVAE compresses initial images into $\times×32$ grids of discrete image tokens, meticulously focusing on learning short-range pixel dependencies. Subsequently, an autoregressive Transformer models the joint distribution over text and image tokens. The overall procedure can be viewed as maximizing the evidence lower bound (ELB) on the joint likelihood of the model distribution over images, captions, and tokens. The evolution of DALL-E into variant models such as DALL-E-2 [100] and DALL-E-3 [9] has unlocked the potential to produce high-resolution images based on detailed and personalized text prompts. DALL-E-2 leverages a CLIP architecture to generate an image embedding induced by the text input and feeds the embedding to a diffusion model to create an image. DALL-E-3 far surpasses its predecessors in creating contextually relevant and visually appealing text embeds by integrating GPT-4, which demonstrates a higher sensitivity to prompt nuances while requiring less user intervention to achieve the desired output. In the realm of ECG analysis, the incorporation of DALL-E as the tokenizer in HeartBEiT [118] facilitates the conversion of ECG plots into discrete tokens, thereby enhancing the overall diagnostic process.
3.1.2 Non-Transformer-based Models.
@@ 4.1 Architecture Design
Both CNNs and MLPs are highly regarded for their ability to effectively model spatial and temporal data. CNN-based architectures have garnered significant attention in self-supervised learning for ECG representation learning, with a notable focus on the utilization of AlexNet [54], EfficientNet [114], ResNet [42] and VGG [108] as common backbones of ECG-FMs. These approaches primarily leverage 1D convolutional operations and pooling operations. 1D convolution aids in learning short-term temporal patterns, such as heartbeat mode, while pooling can reduce data dimensionality while retaining essential features like R-peak information. Examples such as JCDCL [72] and 3KG [34] both utilize a CNN encoder to extract cardiac features, claiming superior efficiency in memory usage and processing speed while delivering competitive performance. Additionally, 1D ResNet is widely used in various pre-training modes, with Adversarial masking [10, 11], METS [63], GCL [107], and SSL-KD [92] employing different 1D ResNet variants as encoders to extract channel-dependent and periodic ECG patterns. On the other hand, MLP-based models are praised for their lightweight design, offering benefits in terms of reduced computational time and cost. For instance, CPC [81] utilizes a simple MLP to encode input sequences, demonstrating high efficiency in forecasting latent representations.
@@ 4.2 Pipeline
3.2.1 Pre-training Paradigms.
@@
Figure 3: Pipeline of ECG-FMs.
Contrastive-based pre-training.
We classify the contrastive-based pre-training strategy as a comprehensive modeling of ECG representations, encompassing augmentation contrast, prediction contrast, and multimodal contrast. As shown in Table 1, augmentation contrast stands as one of the most widely used contrastive frameworks. Most ECG-FMs leverage data augmentation techniques, such as jittering, scaling, and channel masking, to generate different views of an input ECG sample, and then ensure that these augmented views are represented in a similar way in the feature space while pushing representations of different data points far apart. The key lies in constructing and utilizing the self-supervision signals by generating informative positive pairs and filtering out unsuitable negative pairs during augmentation. Due to the periodic nature of the ECG and the significance of the QRS region in containing essential cardiac information, the random heartbeat splitting and joining transformation is also meaningful for ECG augmentation contrast. For instance, sCL-ST [60] splits ECG samples into many snippets and randomizes the concatenation of these snippets to produce augmented views, significantly enhancing the robustness of learned representations across different cardiac diseases. Furthermore, augmentation contrast in the frequency domain is also feasible for ECG data. SSL-KD [92] leverages fast Fourier transform (FFT) to generate a frequency view for ECG tokens.
@@
Reconstruction-based pre-training.

### v3

@@ Introduction
Transformer-based ECG-FMs
The architecture of ECG-FMs increasingly builds on the Transformer, originally developed for NLP tasks [63]. By relying on attention mechanism rather than convolution or recurrence, Transformers can capture long-range dependencies in sequences and scale effectively across domains, driving their adoption in CV [64], finance [65] and healthcare [66].
While the original Transformer uses an encoder–decoder design, recent studies in ECG analysis have explored three main Transformer modes: encoder-only, decoder-only, and encoder–decoder [67]. Encoder-only models, such as ViTs and BERTs, tokenize the input ECG sequence and use self-attention across all tokens to construct contextualized representations, making them well-suited for tasks that require comprehensive signal understanding, such as cardiac disease detection [68]. Decoder-only models, exemplified by GPTs and LLaMAs, generate outputs autoregressively: each new token is predicted based on previously generated ones. This causal structure enables scalable pre-training on massive unlabeled ECG datasets and supports generative tasks such as ECG synthesis [20], interactive assistants that output text-based interpretations of ECG signals [69, 70], and zero-shot inference, i.e., performing tasks without explicit training on them by leveraging representations learned from related tasks [71, 72]. Hybrid encoder–decoder structures combine both approaches: the encoder first produces a global representation of the ECG signal, which the decoder then uses to autoregressively generate outputs conditioned on this representation. Such designs have been applied to diverse ECG tasks, including automated report generation and sequence-to-sequence prediction [73, 74, 70]. Beyond these hybrids, simpler Transformer variants have also been explored for pre-training [75, 76]. These architectural variations illustrate how different Transformer modes offer complementary strengths and underscore the need for continued exploration in ECG analysis. The following section introduces the commonly used models across the three Transformer modes, highlighting the mainstream backbones of ECG-FMs and illustrating their roles in ECG analysis for clinicians and researchers.
Encoder-only ECG-FMs
The encoder-only ECG-FMs typically include three categories: ViT series, BERT series, and CLIP series.
  • ViT series. ViT has emerged as a powerful backbone for ECG foundation models because it represents signals as tokens and applies global self-attention to capture both long-range temporal dynamics and local semantic cues [38]. In ECG tasks, tokens are obtained by segmenting waveforms into time windows or constructing time–lead matrices, with lead and positional embeddings guiding the attention mechanism. Most studies adopt masked autoencoder (MAE) pretraining, where a ViT encoder processes visible tokens and a decoder reconstructs the masked ones, encouraging the encoder to learn transferable spatio-temporal features. Several adaptations highlight this potential. ST-MEM [77] incorporates lead embeddings in a MAE, enabling the ViT encoder to jointly model rhythms and cross-lead relations, improving arrhythmia detection and reduced-lead performance. ECG-MAE [78] extended this with a multi-lead ECG ViT, masking across both time and leads to enhance representations for multi-label cardiac diagnosis. Other works have sought to scale this paradigm to larger datasets or to tailor masking strategies for ECG. For example, MaeFE [79] applies time, lead, and joint masking with ViT encoders to disentangle temporal and spatial features for tasks such as arrhythmia and myocardial infarction classification. MAESSL [80] employs ViT encoders of varying scales (ViT-Base, ViT-Large, and ViT-Huge) pretrained with MAE on clinical 12-lead ECGs, then fine-tunes them for left ventricular systolic dysfunction (LVSD) screening across multi-center cohorts. While most approaches rely solely on ViT encoders, some combine them with other feature extractors. CNN-ViT [81] incorporates convolutional modules for local morphological patterns alongside a ViT encoder for long-range context, using masked reconstruction to unify them. LST [82] further extends ViT-based ECG-FMs to the wearable domain by training on millions of single-lead recordings, confirming their competitiveness with CNN-based encoders.
  • BERT series. BERT [31] introduces masked language modeling (MLM) as a pretraining paradigm that learns contextual embeddings by predicting masked tokens from surrounding sequences. Its success in NLP has inspired numerous domain-specific adaptations, such as BEiT [83] for vision, and BioBERT [84], ClinicalBERT [85], and BioLinkBERT [86] for biomedical text. These models share the common principle of leveraging unlabeled corpora through masked prediction, while adjusting tokenization and pretraining corpora to reflect the properties of each domain. When transferred to ECG analysis, BERT-series encoders are generally applied in two ways: (1) as masked representation learners for ECG signals treated as tokenized sequences, or (2) as text encoders in multimodal frameworks that align ECGs with clinical narratives or medical knowledge.
    In vision-inspired applications, BEiT has been repurposed to capture contextual dependencies in ECG images. HeartBEiT [74] leverages BEiT-style masked image modeling (MIM) to learn contextual embeddings from tokenized ECG plots, leading to improved diagnostic performance across multiple arrhythmia and structural heart disease tasks. Similarly, ECG-GPT [73] employs a BEiT encoder pretrained on ImageNet [87] for automated diagnostic reporting from ECG images, showing that visual token prediction enables the extraction of rich semantic features relevant to clinical interpretation. These studies illustrate how the MLM-to-MIM transfer can extend naturally from text to vision and then to physiological signal images, with the BEiT providing the attention-based modeling capacity. Beyond images, the adaptation of biomedical BERT variants to multimodal ECG–text learning has become a major trend. ClinicalBERT and BioBERT, pretrained on corpora of clinical notes and biomedical literature respectively, have been widely integrated as textual encoders. ETP [88] aligns ECG signals with textual diagnostic descriptions using ClinicalBERT in a CLIP-style contrastive framework, yielding transferable ECG–text embeddings that generalize across datasets. METS [89] demonstrates that even a frozen ClinicalBERT encoder can provide sufficient language supervision for zero-shot ECG classification, highlighting the value of textual grounding in low-label regimes. These approaches underscore how language-pretrained BERT models enrich ECG foundation models by anchoring them to medical semantics. Meanwhile, some efforts have explored treating ECGs themselves as a type of “language". For example, ECGBERT [90] directly tokenizes ECG waveforms into discrete segments and applies MLM pretraining, mirroring the original BERT design but over physiological sequences. This framework demonstrates that ECG-specific masked token prediction can uncover latent temporal and morphological patterns, supporting strong transfer to arrhythmia detection and other diagnostic tasks.
  • CLIP series. CLIP [91] has demonstrated the power of miltimodal contrastive pretraining by aligning images and text within a shared embedding space. Its dual-encoder design, comprising independent Transformer encoders for vision and language, maximizes similarity between paired samples while minimizing it across negatives, enabling broad zero-shot generalization from natural language prompts. Subsequent adaptations, such as MedCLIP [92] and BiomedCLIP [93], have extended this framework to the medical domain through domain-specific corpora and encoders. Building on these developments, CLIP-like approaches have recently been applied to physiological signals like ECGs, where aligning signals with cardiological text facilitates more interpretable and transferable foundation models.
    Early work explored direct adaptation of CLIP to pair ECGs with textual descriptions. CardioGPT [73] adopts a CLIP framewrok by coupling an ECG encoder with a GPT-based text decoder, enabling automatic generation of diagnostic reports from raw signals. Similarly, M3Bind [94] extends this approach by aligning ECGs with shared text embeddings alongside other medical modalities, demonstrating that even the original CLIP architecture can yield clinically meaningful cross-modal retrieval and interpretation. Building on this foundation, subsequent research developed CLIP-like models tailored to ECG-specific challenges such as label diversity, clinical semantics, and limited dataset size. JMCL [95] applies large-scale contrastive pretraining on Japanese ECG datasets with over one hundred diagnostic categories, showing that ECG–text alignment enhances classification performance and facilitates transfer across heterogeneous label spaces. ETP [88] further aligns ECG signals with ClinicalBERT-encoded diagnostic reports, producing embeddings that transferred effectively to arrhythmia classification and generalized across datasets METS [89] further freezes the text encoder to reduce complexity while still enabling zero-shot ECG classification guided by textual prompts, highlighting the efficiency of leveraging pretrained language models for ECG–text alignment.
Decoder-only ECG-FMs
With the rapid progress of LLMs, their scope has extended beyond text-only tasks to encompass multimodal applications. In ECG analysis, both unimodal and multimodal variants (MLLMs) have been investigated, exemplified by models such as Qwen [96] and DeepSeek [30]. As many of these models are grounded in related architectures and share common design principles, we focus on three representative branches of decoder-only models: GPT series, LLaMA series, and DALL-E series.
  • GPT series. GPTs are autoregressive Transformers that generate outputs token by token, conditioning each prediction on previously generated tokens through causal self-attention [97]. Unlike encoder-only architectures that learn bidirectional embeddings, GPTs are decoder-only models designed for sequence generation, making them particularly effective for tasks that require producing structured text or reasoning over sequential inputs. In the context of ECG analysis, GPTs are especially attractive because they serve not only as signal encoders but also as generative interpreters, transforming ECG data into clinically meaningful narratives. When appropriately prompted, they can generate waveform descriptions, diagnostic explanations, or step-by-step reasoning, thereby bridging raw signal analysis with natural language communication in ways well aligned with clinical practice.
    Recent work has proposed multiple strategies for adapting GPT-style architectures to ECG analysis. One approach leverages frozen GPTs as medical knowledge engines. For instance, by prompting pretrained GPT-4 with ECG-derived features or textualized signal representations, Zero-shot RAG [72] demonstrates that GPTs can generate clinically coherent waveform explanations and support diagnostic reasoning without additional task-specific training. Beyond direct prompting, researchers have also explored GPTs as tools for clinical integration and evaluation. SignalGPT [69] incorporates large GPT models into biomedical report-drafting workflows, using their generative capabilities to produce preliminary interpretations that clinicians can refine. ECG-Expert-QA [98] introduces a benchmark for assessing GPT-based models in heart disease diagnosis, systematically testing their ability to answer cardiology-focused questions and highlighting both their promise and current limitations as diagnostic assistants. Meanwhile, ECG-LM [99] further extends this direction by training a BioMedGPT directly on ECG data paired with textual annotations, aiming to develop a unified generative model capable of both understanding and producing ECG-related content.
  • LLaMA series. The LLaMA series extends the decoder-only Transformer architecture with efficiency-oriented refinements such as SwiGLU activations [100] and rotary positional embeddings (RoPE) [101]. These enhancements enable effective handling of long-context sequences and facilitate parameter-efficient fine-tuning, making LLaMAs particularly well suited for ECG analysis, where extended recordings and computational constraints present persistent challenges. In addition, unlike the GPT series, LLaMA has been released as an open-source framework, providing broad access to large-scale generative modeling. This openness has further stimulated rapid progress in healthcare applications, where transparency and reproducibility are especially critical.
    Recent studies have adapted LLaMA models to diverse ECG applications, including signal restoration, diagnostic reasoning, and wearable health monitoring. For example, ECG-LLM [102] addresses noisy or incomplete recordings by casting signal restoration as a generative prediction task, in which tokenized ECG inputs are reconstructed by LLaMA3 into physiologically plausible signals suitable for downstream analysis. Similarly, Zero-shot RAG [72] employs LLaMA2 as the reasoning core, where ECG-derived prompts are enriched with retrieved cardiological knowledge and then processed to produce diagnostic outputs without task-specific fine-tuning. Extending to wearable health applications, CBPM-LLaMA [103] integrates ECG and PPG signals using LLaMA3, guided by textual supervision, to enable accurate non-invasive blood pressure estimation.
  • DALL-E series. DALL-E [104] extends LLMs into text-to-image generation, enabling visual synthesis directly from natural language prompts. Its evolution from the token-based DALL-E-1 to the diffusion-guided DALL-E-2 [105] and DALL-E-3 [106] has progressively improved visual quality and semantic fidelity, with DALL-E-3 demonstrating particularly strong alignment between text prompts and generated images. While originally developed for creative imagery, these models are increasingly being explored in medicine as tools for producing synthetic clinical illustrations and educational materials.
    In ECG-related research, DALL-E has been investigated primarily for visualization and educational support. For instance, DECG [107] systematically evaluates its capacity to generate realistic ECG waveforms and schematic teaching diagrams from textual prompts, which shows that DALL-E-3 could produce visually convincing ECG-like outputs for instructional purposes, but often fails to capture lead-specific morphology with sufficient precision for clinical diagnosis. CDE-MedED [108] demonstrates how DALL-E can support cardiology education by converting textual case descriptions into visual teaching aids, including schematic ECG diagrams and rhythm illustrations, thereby enhancing comprehension and reducing preparation time.
Encoder-decoder ECG-FMs
While encoder-only architectures have proven effective for robust ECG representation learning, and decoder-only models have shown strength in generative reasoning and interpretability, many tasks in ECG analysis demand a combination of both capabilities. Encoder–decoder frameworks can naturally provide this balance: the encoder captures contextualized features from the input sequence, while the decoder generates reconstructions, predictions, or textual outputs conditioned on those representations. In ECG-FMs, encoder-decoder designs appear in two main forms: (1) direct use of the original Transformer architecture with both encoder and decoder blocks, and (2) hybrid systems that integrate encoder-only and decoder-only models to leverage complementary strengths.
  • Simple Transformers. Early studies directly employed the full encoder–decoder Transformer for ECG analysis. For instance, DCL [109] utilizes the encoder to extract signal features while the decoder reconstructs masked segments, with dual contrastive objectives enabling representation transfer to emotion recognition and glucose-level estimation. SK-TF [75] adopts a comparable architecture, pretraining on large ECG datasets before fine-tuning for arrhythmia detection, achieving stronger performance than encoder-only baselines. MTECG [76] further extends this approach by applying a masking strategy across the encoder–decoder pipeline, where the encoder captures contextual representations and the decoder reconstructs masked portions. Collectively, these studies demonstrate that the simple Transformer remains effective for ECG representation learning by combining bidirectional encoding with reconstruction-driven objectives.
  • Hybrid integrations. Beyond the classical design, recent research has advanced toward hybrid architectures that integrate encoder-only and decoder-only models, capitalizing on their complementary strengths for ECG analysis. For example, HeartBEiT [74] combines a BEiT encoder with a DALL-E decoder. The encoder tokenizes ECG images into patch embeddings using masked image modeling, while the decoder reconstructs the missing patches, compelling the encoder to learn rich spatio-temporal representations. ECG-GPT [73] links a BEiT encoder with a GPT-2 [91] decoder, where high-level features extracted by the encoder are passed to the decoder to autoregressively generate textual diagnostic reports. This integration establishes an end-to-end pipeline that translates visual signal inputs into natural language, bridging ECG interpretation with clinical documentation. Building on this framework, other studies have extended the paradigm to multimodal integration. CQA-ESI [70] first employs a CLIP-like encoder to align ECG signals with biomedical text, and then couples it with a GPT-3.5 [110] decoder. The encoder produces multimodal embeddings grounded in both signals and text, while the decoder generates potential waveform details and clinical interpretations by querying external medical knowledge bases. Similarly, ECG-Chat [111] leverages ECG-CoCa as the encoder to integrate ECG and textual representations into a shared space, with GPT-4o [6] serving as the decoder to produce conversational outputs. This design enables interactive diagnostic reasoning and facilitates patient–clinician communication. GEM [112] also employs ECG-CoCa [111] for signal encoding, but explores multiple decoder options, including GPT-4o, SFT-LLaVA, and SFT-PULSE [113]. These decoders generate textual reasoning conditioned on multimodal embeddings, thereby supporting feature-grounded analysis, evidence-driven diagnosis, and clinician-style diagnostic workflows.
Non-Transformer-based ECG-FMs
@@
Both CNNs and MLPs are highly regarded for their ability to effectively model spatial and temporal data. CNN-based architectures have garnered significant attention in self-supervised learning for ECG representation learning, with a notable focus on the utilization of AlexNet [114], EfficientNet [37], ResNet [115] and VGG [36] as common backbones of ECG-FMs. These approaches primarily leverage 1D convolutional operations and pooling operations. 1D convolution aids in learning short-term temporal patterns, such as heartbeat mode, while pooling can reduce data dimensionality while retaining essential features like R-peak information. Examples such as JCDCL [116] and 3KG [117] both utilize a CNN encoder to extract cardiac features, claiming superior efficiency in memory usage and processing speed while delivering competitive performance. Additionally, 1D ResNet is widely used in various pre-training modes, with Advm [118, 119], CPT [16], GCL [120], and SSL-KD [121] employing different 1D ResNet variants as encoders to extract channel-dependent and periodic ECG patterns. On the other hand, MLP-based models are praised for their lightweight design, offering benefits in terms of reduced computational time and cost. For instance, CPC [122] utilizes a simple MLP to encode input sequences, demonstrating high efficiency in forecasting latent representations.
@@
Figure 4: Pipeline of building ECG-FMs.
Pre-training paradigms
@@
Contrastive-based pre-training
classify the contrastive-based pre-training strategy as a comprehensive modeling of ECG representations, encompassing augmentation contrast, prediction contrast, and multimodal contrast. Notably, our review adopts a broad perspective by also including contrastive-like models such as BYOL [123] and VICReg [124], which share the core principle of contrastive pre-training. While these methods are not strictly contrastive in formulation, since they do not rely on explicit negative pairs, they still encourage invariance across different augmented views of the same input. BYOL achieves this by enforcing prediction consistency between an online and a target network, whereas VICReg introduces invariance, variance, and covariance regularization to align augmented representations. Including such methods under a contrastive-like umbrella highlights their conceptual proximity and enables a more unified and systematic discussion of pre-training strategies. As shown in Table 1, augmentation contrast stands as one of the most widely used contrastive frameworks. Most ECG-FMs leverage data augmentation techniques, such as jittering, scaling, and channel masking, to generate different views of an input ECG sample, and then ensure that these augmented views are represented in a similar way in the feature space while pushing representations of different data points far apart. The key lies in constructing and utilizing the self-supervision signals by generating informative positive pairs and filtering out unsuitable negative pairs during augmentation. Due to the periodic nature of the ECG and the significance of the QRS region in containing essential cardiac information, the random heartbeat splitting and joining transformation is also meaningful for ECG augmentation contrast. For instance, sCL-ST [125] splits ECG samples into many snippets and randomizes the concatenation of these snippets to produce augmented views, significantly enhancing the robustness of learned representations across different cardiac diseases. Furthermore, augmentation contrast in the frequency domain is also feasible for ECG data. SSL-KD [121] leverages fast Fourier transform (FFT) to generate a frequency view for ECG tokens.
@@
Reconstruction-based pre-training